OpenStudy (anonymous):

Engineers want to design brakes that will reduce stopping distances by 10%. How would this change the equation relating speed s, friction f, and stoppings=√30fd

OpenStudy (tkhunny):

Is it \(Stopping\;Distance = \sqrt{30\cdot f\cdot d}\), or something else? Maybe \(Speed = \sqrt{30}\cdot Friction \cdot Distance\)